OpenStudy (anonymous):

write (2b)^4 without exponents

OpenStudy (anonymous):

would it be 2b 2b 2b 2b?

OpenStudy (debbieg):

Your expression is A way to write it without exponents, certainly (as long as you make it clear that those are products), but I would rearrange it, and nothing says that you can't combine the numerical product (the 2's). So you have: \((2a)(2a)(2a)(2a) = 2\cdot 2\cdot 2\cdot 2\cdot a\cdot a\cdot a\cdot a\) But you can take the product of those 2's.... \(= 16\cdot a\cdot a\cdot a\cdot a\)